REU 2008: SCHEDULE FOR WEEK ONE

MONDAY MORNING, June 23:

Welcome and opening meeting: 9:30–10:00 in the barn YSP/SESAME training session 10:00–12:00 in the barn Apprentice start-up 10:00–12:00 Ryerson 251 Pizza and coke: 12:15 in the barn

MONDAY–FRIDAY AFTERNOONS: Babai, Farb, May Times listed are maximal: some sessions may end earlier. Babai (Discrete mathematics): 3:30–5:30 MWF, 4:00–5:30 Th; R251 Farb (Topics in geometry): 1:30–3:30 TTh; Ryerson 251 May (Finite spaces): 1:30–3:00 MWF; Eckhardt 133

TUESDAY–FRIDAY mornings, 9:30–12:00 YSP/SESAME training sessions in the barn Babai: Apprentice program in Ryerson 251

TUESDAY 4:00: Meetings to set up mentoring pairings YSP E202, SESAME E203, APPRENTICE E206

Date  YSP/SES  APP  Babai  Farb  May
Mon, June 23 10:00-12:00 10:00-12:00 3:30–5:30  1:30–3:00
Tue, June 24 9:30–12:00 9:30-12:00  1:30–3:30
Wed, June 25 9:30–12:00 9:30-12:00 3:30–5:30  1:30-3:00
Thu, June 26 9:30–12:00 9:30-12:00 4:00–5:30 1:30–3:30
Fri, June 27 9:30–12:00 9:30-12:00 3:30–5:30  1:30-3:00